OpenStudy (wiggler):

Find the horizontal asymptote of (x^2)/sqrt((x^4)+2)

OpenStudy (wiggler):

Horizontal Asymptote of:\[\frac{ x^2 }{ \sqrt{x^4+2} }\]

OpenStudy (anonymous):

try this: devide everything by x^2. You will get: \(1/\sqrt{1+(2/x^4)}\) now just by inspection you find that asymptote is y=1

OpenStudy (wiggler):

Could you please explain dividing the denominator by x^2?

OpenStudy (anonymous):

are you asking how it was done or why it was done?

OpenStudy (wiggler):

how

OpenStudy (anonymous):

this is how: \( \huge\frac{x^2/x^2}{\sqrt{1+2}/x^2}=\frac{x^2/x^2}{\sqrt{1+2/x^4}}=\frac{1}{\sqrt{1+2/x^4}}\)

OpenStudy (wiggler):

okay, thanks :)

OpenStudy (anonymous):

yw